# Evaluating Scoring Bias in LLM-as-a-Judge

### v3

@@ Introduction
We quantify the effects of scoring bias on both scoring accuracy and scoring tendency. Given the golden scores, $S=\{s_{1},\cdots,s_{n}\}$, and the LLM judge’s scores, $S^{\prime}=\{s^{\prime}_{1},\cdots,s^{\prime}_{n}\}$, we quantify scoring accuracy as the correlation between $S$ and $S^{\prime}$. Specifically, we report both Spearman’s correlation coefficient $\rho$ (Spearman 1904) and Pearson’s correlation coefficient $r$ (Pearson 1895). The scoring tendency is quantified by the score distribution:
$N_{k}$ $=\sum_{i=1}^{n}{\mathbb{I}(s_{i}=k)},\quad k\in\{1,2,3,4,5\},$ (3)
$N$ $=(N_{1},N_{2},N_{3},N_{4},N_{5}).$ (4)
Here, $N_{k}$ represents the count of score $k$ in the scoring results, and $N$ denotes the distribution of the score counts.
Experiments
Model Perturbation Strategy BiGGen Bench FLASK MT Bench Vicuna Bench
Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
GPT-4o Baseline - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
Random 0.5963 0.6313 0.7698 0.8124 0.5592 0.5893 0.7420 0.7573 0.7673 0.7838 0.6701 0.7504
Score IDs Letter 0.5641 0.6034 0.7595 0.7994 0.5558 0.5977 0.7341 0.7609 0.7317 0.7454 0.6539 0.7712
Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
DeepSeek-V3 Baseline - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
Random 0.5136 0.5609 0.6547 0.7125 0.4999 0.5470 0.6521 0.6752 0.6178 0.6460 0.5086 0.5482
Score IDs Letter 0.5378 0.5689 0.6803 0.7179 0.5314 0.5755 0.6663 0.6999 0.6489 0.6715 0.4832 0.5657
Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
Qwen3-32B Baseline - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
Random 0.5528 0.5938 0.7404 0.7824 0.5115 0.5456 0.6537 0.6748 0.6554 0.6688 0.6416 0.7252
Score IDs Letter 0.5353 0.5851 0.7280 0.7770 0.5393 0.5724 0.6650 0.6975 0.6795 0.6930 0.6157 0.6808
Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
Qwen3-8B Baseline - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
Random 0.4869 0.5025 0.6091 0.5950 0.4778 0.5004 0.6061 0.6216 0.5949 0.5983 0.5135 0.6053
Score IDs Letter 0.4853 0.5117 0.6410 0.6665 0.4873 0.5244 0.6334 0.6586 0.5985 0.6126 0.5812 0.6890
Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
Mistral-24B Baseline - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
Random 0.5053 0.5450 0.6569 0.6841 0.4756 0.5042 0.6212 0.6338 0.5603 0.5602 0.5461 0.6418
Score IDs Letter 0.5114 0.5632 0.6685 0.7084 0.4880 0.5339 0.6311 0.6594 0.5375 0.5485 0.5558 0.6608
Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
Table 3: Correlation between golden scores and judge model scores under various prompt perturbations. For each judge model, the first row (Baseline) presents the metrics for the original, unperturbed prompt. The subsequent rows show the results under different types of perturbations applied to the rubrics order and score IDs. The degradation in correlation scores compared to the baseline reflects the model’s scoring bias, indicating its lack of robustness to these perturbations. Model names are abbreviated: DeepSeek-V3 denotes DeepSeek-V3-67B and Mistral-24B denotes Mistral-Small-24B-Instruct. We report Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient.
Setup
@@
Judge Models. Previous work (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability (Ye et al. 2024a). After preliminary experiments, we found that scoring bias is a common phenomenon in judge models based on various LLMs. Furthermore, we aim to explore the impact of factors such as model type and size on scoring bias. Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models. They are GPT-4o, DeepSeek-V3-671B (Liu et al. 2024a), Qwen3-32B, Qwen3-8B (Yang et al. 2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025). These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al. 2023; Ye et al. 2024a).
@@
Judge Prompt. The prompt we use for scoring-based judgment, adapted from (Kim et al. 2024c), is shown in Figure 3. The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
@@
Main Results
Figure 5: Spearman’s correlation between human annotator scores and judge model scores on the FLASK dataset, under perturbations applied to the reference answer scores.
@@
The existence of scoring bias is proved in all judge models. Table 3 and Figure 5 show the correlation between judge model scores and golden scores under all perturbation methods. For all judge models we evaluated, including advanced models such as GPT-4o, we observed fluctuations in the scoring results under different types of perturbations. This indicates that scoring bias significantly impacts the robustness of LLM-as-a-Judge. Furthermore, different judge models exhibit varying sensitivity to scoring bias:
@@
(1) More powerful judge models tend to be more robust to scoring bias. Consistent with existing consensus (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring. As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains within 0.03, except for the case involving a reference answer with a score of 5. A comparison of the results for Qwen3-32B and Qwen3-8B shows that judge models with larger parameter sizes demonstrate greater robustness in scoring. In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubric order bias and score ID bias. With reference answer score bias, the maximum fluctuation reaches nearly 0.2. However, the changes in the correlation coefficient of Qwen3-32B under biases are generally within 0.03. Therefore, more capable judge models should be prioritized in scoring evaluations.
@@
(3) Judge models have their own scoring tendencies. Figure 6 presents the scoring tendencies of different judge models under various scoring biases on the FLASK dataset in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to assign a score of 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for a score of 4, while Qwen3-32B favors a score of 5. An examination of Figure 6 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 6: Score distributions of LLM judges under various prompt perturbations on the FLASK dataset. The horizontal axis represents the assigned score (where -1 indicates a failed judgment), and the vertical axis lists the different perturbation methods applied. For reference, the top row in each heatmap (Human) shows the distribution of human scores. The second row (None) shows the model’s score distribution for the original, unperturbed prompt.
Methods for Scoring Bias Mitigation
For advanced or large size judge models, applying scoring prompt that deviate from conventional preferences is worth considering.
@@
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubric order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models. Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention. Effective strategies include introducing structured, unconventional formats, such as presenting scoring rubrics in a descending order or using non-numeric identifiers like letter grades (A, B, C). However, the key is to maintain clarity and logical coherence. Therefore, chaotic perturbations—such as a fully randomized rubric order or non-hierarchical identifiers—should be avoided. Moreover, inspired by previous work (Liu et al. 2024c; Park et al. 2024), perturbations applied to scoring prompt also show potential as debiasing strategies. For example, effective approaches include implementing multiple variations of the scoring rubric order or IDs and then aggregating the results, or adversarially training the models on samples enhanced with scoring bias perturbations.
@@
Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 5 and Figure 6. The experimental results indicate that, compared to the score rubric order bias and the score ID bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and more regular. Figure 5 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 5 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
@@
Figure 6 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer is between 1 and 4, these models tend to assign scores similar to that of the reference answer. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
@@
The above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
Limitations and Future Work

### v4

@@ 3.4 Metrics
4 Experiments
4.1 Setup
@@
Judge Models. Previous work [7, 18, 21] has reached a consensus that judge models should be stronger models to ensure evaluation reliability [47]. After preliminary experiments, we found that scoring bias is a common phenomenon in judge models based on various LLMs. Furthermore, we aim to explore the impact of factors such as model type and size on scoring bias. Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models. They are GPT-4o, DeepSeek-V3-671B [20], Qwen3-32B, Qwen3-8B [46], and Mistral-Small-24B-Instruct-2501 [24]. These judge models we selected differ from the models used to generate responses in the evaluation datasets, thus mitigating potential self-enhancement bias [47, 51].
@@
Judge Prompt. The prompt we use for scoring-based judgment, adapted from [13], is shown in Figure 3. The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
@@
4.2 Main Results
Table 3: Stability metrics FP (%) and MAD, comparing baseline scores with scores under various prompt perturbations. A higher metric indicates a greater difference from the baseline score. The perturbation method Ref-5 was only applied to BiGGen Bench and FLASK, as the other two datasets do not provide reference answers.
Model Perturbation BiGGen Bench FLASK MT Bench Vicuna Bench
FR MAD FR MAD FR MAD FR MAD
GPT-4o Descending-Numeric 23.63 0.2885 23.44 0.2769 23.13 0.2594 20.00 0.2094
Random-Numeric 22.55 0.2651 24.94 0.3023 28.44 0.3375 19.38 0.1938
Letter-Grades 18.17 0.2122 18.24 0.2259 19.69 0.2188 19.38 0.2031
Roman-Numerals 16.51 0.1892 15.94 0.2014 23.13 0.2688 15.63 0.1625
Ref-5 45.54 0.5604 45.58 0.7166 - - - -
DeepSeek-V3-671B Descending-Numeric 24.21 0.2867 21.89 0.2734 25.00 0.2969 27.50 0.2906
Random-Numeric 30.86 0.3565 32.83 0.4123 32.19 0.3875 39.06 0.4250
Letter-Grades 19.86 0.2349 18.14 0.2374 17.81 0.1969 26.25 0.2844
Roman-Numerals 19.53 0.2338 17.89 0.2169 20.00 0.2344 25.00 0.2625
Ref-5 38.99 0.4863 45.13 0.6742 - - - -
Qwen3-32B Descending-Numeric 28.56 0.3281 25.54 0.2989 29.38 0.3656 34.06 0.3688
Random-Numeric 26.12 0.2957 29.64 0.3448 33.44 0.3938 27.81 0.2906
Letter-Grades 25.11 0.2770 24.34 0.2909 28.44 0.3313 29.06 0.3000
Roman-Numerals 28.92 0.3259 28.39 0.3273 51.88 0.6875 30.94 0.3156
Ref-5 37.59 0.7674 44.58 0.6512 - - - -
Qwen3-8B Descending-Numeric 46.22 0.5296 32.23 0.3778 40.31 0.4594 44.06 0.4750
Random-Numeric 34.93 0.3989 36.83 0.4363 35.31 0.4125 34.38 0.3750
Letter-Grades 30.54 0.3493 28.39 0.3228 36.88 0.4063 28.44 0.2938
Roman-Numerals 32.99 0.3827 34.38 0.4058 40.31 0.4344 30.94 0.3156
Ref-5 36.48 0.4367 40.53 0.5712 - - - -
Mistral-Small-24B-Instruct Descending-Numeric 31.04 0.3421 27.39 0.2904 32.81 0.3406 25.63 0.2719
Random-Numeric 28.02 0.3083 33.03 0.3603 29.38 0.3066 22.81 0.2406
Letter-Grades 25.36 0.2734 25.54 0.2809 26.88 0.2969 23.75 0.2438
Roman-Numerals 21.83 0.2349 23.24 0.2549 21.56 0.2281 16.56 0.1688
Ref-5 35.40 0.4065 48.58 0.6247 - - - -
Table 4: Accuracy metrics Spearman’s ($\rho$) and Pearson’s ($r$) correlation coefficient, showing the correlation between judge scores under perturbations and golden scores. Each judge model’s first row serves as the unperturbed baseline. Results are colored green (better than baseline) or red (worse).
Model Perturbation Human Score GPT-4.1 Score
BiGGen Bench FLASK MT Bench Vicuna Bench
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
GPT-4o - 0.5669 0.6147 0.5529 0.5915 0.7578 0.7724 0.6862 0.7696
[…]
DeepSeek-V3-671B - 0.5340 0.5823 0.5193 0.5746 0.6323 0.6593 0.4411 0.6119
[…]
Qwen3-32B - 0.5418 0.5848 0.5323 0.5695 0.6959 0.7053 0.6419 0.7188
[…]
Qwen3-8B - 0.5087 0.5419 0.5035 0.5309 0.6238 0.6342 0.6639 0.7284
[…]
Mistral-Small-24B-Instruct - 0.5067 0.5543 0.4948 0.5392 0.5943 0.5762 0.6116 0.6917
[…]
@@ 4.2 Main Results
The existence of scoring bias is proved in all judge models. Table 3 shows the difference between baseline model scores and perturbed model scores, Table 4 and Figure 6 show the correlation between judge model scores and golden scores under all perturbation methods. For all judge models we evaluated, including advanced models such as GPT-4o, we observed fluctuations in the scoring results under different types of perturbations. This indicates that scoring bias significantly impacts the robustness of LLM-as-a-Judge. Furthermore, different judge models exhibit varying sensitivity to scoring bias:
@@
(1) More powerful judge models tend to be more robust to scoring bias. Consistent with existing consensus [7, 18, 21], our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability in scoring. As shown in Table 3, GPT-4o has FRs less than 25% and MADs less than 0.3 under perturbation in most cases, except for the case involving a full-marked reference answer. Althogh Qwen3-8B, has higher FRs and MADs than GPT-4o and the larger version Qwen3-32B, especially reaches 46.22% FR and 0.5296 MAD on BiGGen Bench under descending-numeric perturbation, which means that nearly half of the scores are corrupted. Therefore, more capable judge models should be prioritized in scoring evaluations.
@@
(3) Judge models have their own scoring tendencies. Figure 5 presents the scoring tendencies of different judge models under various scoring biases in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to assign a score of 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. For example, GPT-4o demonstrates a preference for a score of 4. An examination of Figure 5 reveals that score rubric order bias and score ID bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
(a) BiGGen Bench
(b) FLASK
Figure 5: Score tendency of LLM judges on BiGGen Bench and FLASK, plotting score distributions (x-axis) against perturbation methods (y-axis). The “Human” (top) and “None” (second) rows serve as baselines, showing human scores and the model’s unperturbed scores, respectively.
4.3 Methods for Scoring Bias Mitigation
4.3.1 (1) Leveraging Model Scale as a Debiasing Strategy.
Our findings show a strong correlation between model scale and bias resilience. As shown in Table 3, state-of-the-art models like GPT-4o consistently demonstrate high stability, with FRs often remaining below 25%. In stark contrast, smaller models are highly vulnerable. Qwen3-8B, for example, suffers a 46.22% FR on BiGGen Bench under a simple rubric reordering. The larger Qwen3-32B is far more stable (28.56% FR) under the same perturbation, confirming the link to model scale. This trend extends to accuracy, as shown in Table 4. Perturbations often improve GPT-4o’s correlations with human scores, whereas the same perturbations frequently degrade the accuracy of smaller models like Qwen3-8B and Mistral-Small. Therefore, for high-stakes evaluations, selecting a powerful model like GPT-4o is a primary strategy for ensuring robust and reliable results, justifying potential trade-offs in cost or latency.
4.3.2 (2) Debiasing via Intentional Deviation from Convention.
@@ 4.3 Methods for Scoring Bias Mitigation
In most cases of scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs, and the score rubric order typically follows a sequential order (1 to 5). These conventional choices align with human cognitive preferences in prompt design. However, our experimental results (Table 4) indicate that such prompts do not necessarily produce the most accurate scoring. For example, the GPT-4o judge improves when using Roman numerals. Besides, using letter grades as scoring IDs improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubric order bias, adopting a descending order improves the scoring accuracy of DeepSeek-V3 and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubric order tends to negatively impact the scoring accuracy of all judge models. Based on our analysis, we propose a core design principle for scoring prompts: intentional yet meaningful deviation from convention. Effective strategies include structured, unconventional formats, such as presenting rubrics in descending order or using non-numeric identifiers like letter grades (A, B, C). However, the key is to maintain clarity. Therefore, chaotic perturbations—such as a fully randomized rubric order—should be avoided as they harm logical coherence. Moreover, inspired by previous work [23, 29], perturbations applied to scoring prompt also show potential as debiasing strategies.
@@
4.3.3 (3) Using full-mark reference answers can stabilize scoring preferences and achieve optimal accuracy.
The impact of reference answer score bias on scoring accuracy and tendency is shown in Figure 6 and Figure 5. Our results indicate this bias’s influence on robustness is more pronounced and regular than that of score rubric order or ID bias.
As shown in Figure 6, introducing reference answers generally improves scoring accuracy for models like GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct. Notably, across both BiGGen Bench and FLASK, introducing a full-marked (score 5) reference answer consistently yields top-tier scoring accuracy, significantly outperforming other scores and the baseline.
@@
Figure 5 reveals a significant impact on scoring tendency, especially for DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B. When the reference score is between 1 and 4, these models tend to assign similar scores, indicating a strong pull from the reference. Interestingly, when a score 5 reference answer is introduced, the judge’s scoring behavior becomes more rational.
@@
In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, when incorporating reference answers into a scoring-based judge prompt, a full-mark reference answer is the best choice. Conceptually, this aligns with metrics like BLEU and ROUGE, which also rely on golden answers.
(a) BiGGen Bench
(b) FLASK
Figure 6: Accuracy metric Spearman’s correlation between human scores and judge model scores on BiGGen Bench and FLASK dataset, under perturbations applied to the reference answer scores.
5 Limitations and Future Work